OpenStudy (anonymous):

What is the 5th term of an arithmetic sequence if t3 = 10 and t7 = 26

OpenStudy (cwrw238):

t7 - t3 = 16 so common deference is 16/4 = 4 so fifth term = t3 + 2*4 = 10 + 8 = 18

OpenStudy (cwrw238):

* common difference

OpenStudy (cwrw238):

do u follow that ?

OpenStudy (anonymous):

thank you so muchhhhhhhhh

OpenStudy (cwrw238):

yw